OpenStudy (angel123):

PLZZ HELPPP!!!To get out of jail free in the board game MONOPOLY, you have to roll doubles with a pair of standard dice. If you fail in three attempts, you have to pay a fine to get out of jail. Determine the probability of getting out of jail without paying a fine.

jimthompson5910 (jim_thompson5910):

look at a dice chart like this http://www.bkgm.com/articles/GOL/gif/hank2x.gif how many cases do doubles show up? A "double" is where you roll the same number at the same time (eg: rolling two '6's)

OpenStudy (angel123):

6/36

jimthompson5910 (jim_thompson5910):

which reduces to 1/6. Agreed?

OpenStudy (angel123):

Yes

jimthompson5910 (jim_thompson5910):

The probability of rolling doubles is 1/6 What is the probability of not rolling doubles?

OpenStudy (angel123):

30/36 - 5/6

jimthompson5910 (jim_thompson5910):

1 - (1/6) = 6/6 - 1/6 = 5/6 I think that's what you mean to say?

OpenStudy (angel123):

yea

jimthompson5910 (jim_thompson5910):

the probability of not rolling doubles is 5/6 multiply that with itself 3 times to get (5/6)*(5/6)*(5/6) = ???

OpenStudy (angel123):

125/216

jimthompson5910 (jim_thompson5910):

that is the probability of rolling 3 non-doubles it is the probability of having to pay the fine

jimthompson5910 (jim_thompson5910):

subtract 125/216 from 1 to get the final answer

jimthompson5910 (jim_thompson5910):

1 - (125/216) = ???

OpenStudy (angel123):

OMG i got it thanks alot so it would be 91/216

OpenStudy (angel123):

right?

jimthompson5910 (jim_thompson5910):

91/216 is correct

OpenStudy (angel123):

okay thanks alott!!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (vuriffy):

@jim_thompson5910 I should of got on here for help with probability last year, you would of helped so much. Haha.